if a varies directly as the cube of c, and a = 18 when c = 30, find a when c = 6.
Answer by ewatrrr(24785)


Hi
a = k(c^3) |where k is the constant of the direct variation
18 = k(30^3)
18/27000 = .00067
a = .00067(6^3)= .1447